### v1

@@ 3.1 Details of numerical simulations
$u_{0}(x)=\begin{cases}0.9\ K\quad&\text{for }\lvert x\rvert\leq R_{u}\\ 0\ \quad&\text{for }\lvert x\rvert>R_{u}\end{cases}\qquad i_{0}(x)=0\qquad v_{0}(x)=\begin{cases}V_{0}\quad&\text{for }\lvert x\rvert\leq R_{v}\\ 0\quad&\text{for }\lvert x\rvert>R_{v}\end{cases}$ (3.1)
with $V_{0}=2.67\times 10^{4}$ in one spatial dimension and $V_{0}=(2.67\times 10^{3})^{2}$ in two spatial dimensions; these values are obtained by appropriately rescaling the one used in Ref. [56], which comes from the number of viral particles used in Ref. [30]. The same initial equations are also employed for the agent-based models.
@@
The viral load released by the death of infected cells $\alpha$ depends highly on the type of virus and ranges from the value 157157\;viruses/cells estimated in Ref. [71] to the value 35003500\;viruses/cells of Ref. [13]; this wide variation is among the motivations that led us to consider different values of $\alpha$. The value of Ref. [71] however appears too small to show an effective infection in our settings, hence in our simulation we use a minimum value of $\alpha$ equal to 580580\;viruses/cells. The clearance rate of the virus is also a parameter that could significantly vary in different situations: in Ref. [44], it has been estimated to $1.67\times 10^{-1}\;$h-1 in vitro, but it is reasonable to assume it to be higher in different settings, due for example to the immune response in vivo or other factors that reduce the viral lifespan. We therefore also consider a higher value of 1.001.00\;h-1. We finally remark that an adequate comparison requires the adaptation of the diffusion coefficient to keep the propagation speed constant in the different models (as listed in Table 1 and further explained in Appendix A.1).
All the simulations have been performed in Matlab 2021b with the techniques described in Ref. [46]. For the individual-based models, we used a temporal step $\tau=0.02\;$h and a spatial step of $\delta=0.1\;$mm both for the one-dimensional and the two-dimensional simulations. Since we only need to keep track of the collective fate of cells in the same lattice point, we used the built-in Matlab functions binornd and mnrnd, which compute random arrays according to binomial and multinomial distributions. In order to allow reproducibility, a random seed has been set at the beginning of each new simulation, ranging from one to five (except in the case of Fig. 11, which shows the results of one hundred simulations). In the figures representing a single simulation, only the one with random seed equal to one is shown: we remark that we performed at least five different simulations also in these cases and the resulting stochastic behaviour arises in all of them in a qualitatively similar way (although the exact quantitative features are clearly different).
Eqs. (U3) and (U2) have been solved with a finite difference scheme explicit in time, using the discretisations $\Delta x=0.1$, $\Delta t=10^{-3}$. We used a forward upwind scheme for the nonlinear diffusion of Eqs. (P3) and (P2), following Ref. [36]; the discretisations are again $\Delta x=0.1$, $\Delta t=10^{-3}$. We remark that the simulations of the electronic supplementary material S4 require a refined discretisation $\Delta x=0.01$, $\Delta t=10^{-4}$ to reduce the diffusion due to the numerical method. In the two-dimensional case we exploited the radial symmetry of the equations.
@@ 3.2 Partial treatment success: travelling waves
Figure 1: (a) Theoretical speed of the travelling wave of viral infection invading an uninfected tumour at carrying capacity, computed following the procedure of Ref. [4] (as explained in the Appendix A). The parameters employed are the ones given in Table 1, with three values of $D_{v}$; $q_{v}$ and $\alpha$ are varied so that the value of $\tilde{\beta}$ is constantly equal to $1.02\times 10^{-1}\;$h-1. The dashed cyan line shows the speed of the infection wave obtained for Eq. (U2) with the same parameters. (b) Numerical solutions of Eq. (U3) compared with the position of the theoretical front (shown as a dashed red line). The parameters employed are the ones given in Table 1, with $\alpha=580\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1; we also set $R_{u}=L$ to focus on the infection inside an established tumour.
Let us first focus on the dynamics that do not present persistent wide oscillations. The spatially homogeneous steady-state of the tumour is independent of $D_{v}$ and depends on $\alpha$ and $q$ only through $\tilde{\beta}$. This observation is already enough to provide useful insights on how to improve the overall treatment outcome, which remain valid whenever the infection spreads in the whole tumour. Indeed, increasing $\tilde{\beta}$ in any way and decreasing $q$ appears beneficial to the treatment. Even for parameter sets that maintain the equilibria unchanged, it is reasonable to expect differences in the spatial viral spread, which could significantly affect the treatment.
@@
The speed of travelling waves solutions for PDEs with spatial diffusion can be computed using linearisation techniques. In the case of Eq. (U2), infected cells invade regions in which uninfected cells are at carrying capacity at speed $2\sqrt{D(\beta-q)}$ and slow down as the uninfected cell density decreases [46]; the same result can be also obtain in a rigorous way for a similar system of PDEs [16]. In the case of Eq. (U3), the minimal speed can be found at the intersection of two appropriate manifolds [4]: although it is hard to derive an explicit expression, we can easily rely on numerical computations (see Appendix A). Fig. 1 shows the behaviour of the minimal speed for different parameter values. For the sake of simplicity, we restrict our attention to parameter combinations that maintain the ratio $\alpha/q_{v}$ constant: while this is not a biological constraint, it allows to focus our attention to the specific differences between the modelling approaches in comparable settings. As $q_{v}$ grows, the speed of the infection wave converges to the value $2\sqrt{D(\beta-q)}$, meaning that viral dynamics are irrelevant because the virus is cleared very quickly. Instead, the case $q_{v}=0$ corresponds to $\alpha=0$, which means that viral particles only diffuse without any production or decay; travelling waves are not possible in this hypotetical situation, even though this state has no biological meaning. It is also interesting to observe that for very low values of $q_{v}$ the speed is extremely low: in this case, viral dynamics are mostly dictated by spatial diffusion and the effects of the reaction part of the equation is negligible. In the intermediate cases, the speed is either higher or lower than $2\sqrt{D(\beta-q)}$ depending on whether $D_{v}$ is higher or lower than $D$. In the former case, there is a specific value of $q_{v}$ that maximises the speed.
@@
Figure 3: Comparison in one spatial dimension between numerical simulations of the discrete model with undirected cell movement (solid lines), the numerical solution of Eq. (U2) (dashed lines) and the numerical solution of Eq. (U3) (dotted black lines) at three different times, with the parameters given in Table 1, $\alpha=3500\;$viruses/cells and $q_{v}=1\;$h-1. All the graphical elements have the same meaning as Fig. 3.
Let us now describe how the different speed of the infection affects the treatment in comparison to the previous approaches. We start by considering the lower values of $\alpha$ and $q_{v}$ from Table 1, as in this case we expect a more significant difference. Fig. 2, along with the video accompanying it (see electronic supplementary material S2), shows an excellent quantitative agreement between numerical solutions of the system of PDEs (U3) and the average over five numerical simulations of the corresponding agent-based model in one spatial dimension. Furthermore, the quasi-steady approximation of Eq. (2.3) holds most of the time, even in the agent-based model (despite some very small stochastic fluctuations), as it is evident from Fig. 2a. Nevertheless, the system’s transient behaviour significantly differs from the solution of Eq. (U2). Fig. 2a shows that the low viral decay allows the viral infection to reach the boundary of the tumour faster than in the case of cell-to-cell infection, in line with the higher speed predicted by theoretical results (see the vertical dash-dotted lines); as a consequence, uninfected cells start to decrease their numbers earlier (Fig. 2b). This difference does not significantly affect the system’s overall behaviour: in all the cases, the travelling wave of uninfected cells invades the surrounding area at the speed $2\sqrt{Dp}$ (vertical dashed lines in Fig. 2b-c) and cell densities at the centre of the tumour converge with damped oscillations to the equilibrium of the corresponding ODE (horizontal solid black lines in Fig. 2), given by Eq. (2.6)).
@@
In the full system, we observe a lower cell density at the front, which can be explained by the higher infiltration of the infection: the theoretical approximation predicts this decrease, but significantly underestimates the value; this disagreement is probably caused by the fact that the infection is not invading a spatially homogeneous uninfected population, which is the situation considered in the Appendix. Indeed, such a wave would have a longer spatial decay towards infinity, but this is not possible in the case of Fig. 2c because the uninfected population is more localised; as a consequence, the desired speed can only be obtained if the uninfected peak is a bit higher. In other words, nonlinear effects here play a more important role than in other settings.
Let us now consider higher values of $\alpha$ and $q_{v}$, keeping their ratio approximately as in the previous case. Fig. 3, along with the video accompanying it (see electronic supplementary material S3), shows again an excellent quantitative agreement between numerical solutions of the system of PDEs (U3) and the average over five numerical simulations of the corresponding agent-based model in one spatial dimension. Furthermore, the initial invasion of infected cells shows a larger degree of similarity to the one obtained by the two-equation model in Eq. (U2) than in the previous case, as expected by the value of the invasion speed; nevertheless, the shape of the infected cells shows some differences. After an initial transient time, the uninfected cell density is almost indistinguishable from the one resulting from Eq.  (U2) (Fig. 3b-c). At later times, Fig. 3c shows the formation of travelling waves remarkably similar to the solution of Eq.  (U2), with almost the same density of uninfected cells at the front: the theoretical value is a much better approximation than in the previous case and correctly predicts the front to be slightly lower than for the two equations model.
@@
Let us first consider the lower values of $\alpha$ and $q_{v}$. The behaviour of numerical simulations is very similar to the one observed in the previous case. Let us recall that, in the case of Eq. (P2), central infections remain localised in the initial position, as described in Ref. [46] (see also Subsection 3.3); this is in sharp contrast with the outcome shown in Fig. 4. Indeed, in the present situation, viral particles can reach the outer region of the tumour without any obstacle and the constraints of cell movement do not stop the propagation of the infection. The similarity of this situation to the case of undirected cell movement suggests that it is more appropriate to compare it to Eq. (U2) than to Eq. (P2). Fig. 4a, along with the video accompanying it (see electronic supplementary material S2), shows that the spread of the infection is extremely similar with respect to the previous case, meaning that the viral diffusion affects the dynamics more significantly than cellular movement; small differences can only be observed by a close comparison of the PDE models, which however are totally unnoticeable in the agent-based model. The only relevant differences between undirected movement and standard diffusion appear to be the maximum value reached by uninfected cell density, which is higher in the former situation, and the profile of the final invasion front, which, in the latter case, is compactly supported and propagates to the surrounding tissues slightly faster. Overall, we can conclude that spatial viral dynamics have a more substantial effect on the evolution of the system than cellular spatial dynamics.
Fig. 4c, along with the video accompanying it (see electronic supplementary material S3), shows that also for high $\alpha$ and $q_{v}$ the outcome is very similar to the case of undirected cell movement, which confirms our previous observation that the infective dynamics are affected mainly by viral spatial spread.
@@ 3.3 Possible treatment failure: low viral diffusion
Lower viral diffusion clearly implies a slower invasion speed of infection, as also shown in Fig. 1. The movement of cells is still enough to spread the infection all over the tumour, as shown in Fig. 5a-b. It is also interesting to observe that lower values of $\alpha$ and $q_{v}$ are now associated with a slower speed, differently from the case of faster viral diffusion: indeed, viral particles with a shorter lifespan cannot effectively propagate for long distances. Fig. 5a also shows that some stochastic events slow down the infection to the left of the domain, even though the results are averaged over five simulations; it appears therefore that stochasticity becomes more relevant for this less effective infection, although the overall behaviour is not affected significantly. Fig. 5b shows that for higher values of $\alpha$ and $q_{v}$ the behaviour is almost indistinguishable to the one of Eq. (U2): this is indeed the parameter range in which the quasi-steady assumption of the virus is most accurate. In both cases, the theoretical predictions of the peak of uninfected cells is very accurate for the PDE and still constitutes a good approximation for the agent-based model.
Figure 6: Numerical simulation of the discrete model with pressure-driven movement in two spatial dimensions at three different times. The parameters employed are the ones given in Table 1, with the exception of the diffusion coefficient of viral particles $D_{v}$ (which is set to $10^{-5}\;$mm2/h), the death rate of infected cells $q$ (which is set to $8.33\times 10^{-3}\;$h-1, i.e. one-fifth of the reference values) and the initial radius of viral infection $R_{v}$ (which is set equal to $R_{u}$); furthermore, we use the values $\alpha=3500\;$viruses/cells and $q_{v}=1\;$h-1. The dashed cyan circles in the top panels represent the expected positions of the uninfected invasion fronts in the absence of treatment, travelling at speed $\sqrt{D_{u}p/2}$. The dashed red circles in the bottom panels represents the front of the infected cells given by the numerical solution of Eq. (P3).
@@
As the spatial spread of viral particles becomes less relevant, we expect to recover the main features of Eq. (P2). First of all, a central infection remains mostly confined at the centre of the tumour for both parameter sets that we consider (only shown in the electronic supplementary material S4): the central infection causes the total cell density to drop, while external uninfected cells proliferate; since cells cannot move towards an area with higher cell density and viral diffusion is negligible, the outer cells are never going to be infected and the tumour keeps expanding in the same way it would do in the absence of treatment, almost identically to the situation described in Ref. [46].
We therefore focus on initial infections spread in the whole tumour, namely $R_{v}=R_{u}$, which gives rise to the formation of travelling waves for Eq. (P2) [46]. For the sake of brevity, we restrict our attention to higher values of $\alpha$ and $q_{v}$, which facilitate the spread of the infection for the continuous model; we remark that both cases are shown in electronic supplementary material S5. Fig. 5c shows that the same kind of travelling waves is also obtained as a solution of Eq. (P3) and the two solutions superimpose almost exactly; the peak of uninfected cells is higher than the one of the solutions of Eqs. (U2) and (U3), in line with the less effective spread of the infection. Note that the agent-based model again shows that the infection is confined to the centre of the tumour, in line with the case that neglects explicit viral dynamics. It is important to stress that, for the current model, demographic stochasticity plays a much more important role than in previous models: any growth above the average of uninfected cells stops the movement of infected cells and hence cannot be compensated at later times by other processes.
@@
Reasonable increases of the effective infection rate $\tilde{\beta}$ do not lead to a better infiltration of the infection, as this causes a decrease in central cell density and creates the need for infected cells to move against a pressure gradient. Furthermore, and as we will show shortly, oscillations emerge in some parameter ranges associated with an effective infection, but are suppressed at low values of viral diffusion. On the other hand, building upon the results of Ref. [46], we expect that decreasing the death rate of infected cells $q$ leads to a better outcome.

### v2

@@ 3.1 Details of numerical simulations
$u_{0}(x)=\begin{cases}0.9\ K\qquad&\text{for }\lvert x\rvert\leq R_{u}\\ 0\ \qquad&\text{for }\lvert x\rvert>R_{u}\end{cases}\qquad i_{0}(x)=0\qquad v_{0}(x)=\begin{cases}V_{0}\qquad&\text{for }\lvert x\rvert\leq R_{v}\\ 0\qquad&\text{for }\lvert x\rvert>R_{v}\end{cases}$ (3.1)
with $V_{0}=2.67\times 10^{4}$ in one spatial dimension and $V_{0}=(2.67\times 10^{3})^{2}$ in two spatial dimensions; these values are obtained by appropriately rescaling the one used in Ref. [pooladvand21], which comes from the number of viral particles used in Ref. [kim06]. The same initial conditions are also employed for the agent-based models.
@@
The viral load released by the death of infected cells $\alpha$ depends highly on the type of virus and ranges from the value 157157\;viruses/cells estimated in Ref. [workenhe14] to the value 35003500\;viruses/cells of Ref. [chen01]; this wide variation is among the motivations that led us to consider different values of $\alpha$. The value of Ref. [workenhe14] however appears too small to show an effective infection in our settings, hence in our simulation we use a minimum value of $\alpha$ equal to 580580\;viruses/cells. The clearance rate of the virus is also a parameter that could significantly vary in different situations: in Ref. [mok09], it has been estimated to $1.67\times 10^{-1}\;$h-1 in vitro, but it is reasonable to assume it to be higher in different settings, due for example to the immune response in vivo or other factors that reduce the viral lifespan. We therefore also consider a higher value of 1.001.00\;h-1. We finally remark that an adequate comparison requires the adaptation of the diffusion coefficient to keep the propagation speed constant in the different models (as listed in Table 1 and further explained in Appendix A.1).
All the simulations have been performed in Matlab 2021b with the techniques described in Ref. [morselli23]. For the individual-based models, we used a temporal step $\tau=0.02\;$h and a spatial step of $\delta=0.1\;$mm both for the one-dimensional and the two-dimensional simulations. Since we only need to keep track of the collective fate of cells in the same lattice point, we used the built-in Matlab functions binornd and mnrnd, which compute random arrays according to binomial and multinomial distributions. In order to allow reproducibility, a random seed has been set at the beginning of each new simulation, ranging from one to five (except in the case of Fig. 10, which shows the results of one hundred simulations). In the figures representing a single simulation, only the one with random seed equal to one is shown: we remark that we performed at least five different simulations also in these cases and the resulting stochastic behaviour arises in all of them in a qualitatively similar way (although the exact quantitative features are clearly different).
Eqs. (U3) and (U2) have been solved with a finite difference scheme explicit in time, using the discretisations $\Delta x=0.1$, $\Delta t=10^{-3}$. We used a forward upwind scheme for the nonlinear diffusion of Eqs. (P3) and (P2), following Ref. [leveque07]; the discretisations are again $\Delta x=0.1$, $\Delta t=10^{-3}$. We remark that the simulations of the electronic supplementary material S4 require a refined discretisation $\Delta x=0.01$, $\Delta t=10^{-4}$ to reduce the diffusion due to the numerical method. In the two-dimensional case we exploited the radial symmetry of the equations.
@@ 3.2 Partial treatment success: travelling waves
Figure 1: (a) Theoretical speed of the travelling wave of viral infection invading an uninfected tumour at carrying capacity, computed following the procedure of Ref. [baabdulla23] (as explained in the Appendix A). The parameters employed are the ones given in Table 1, with three values of $D_{v}$; $q_{v}$ and $\alpha$ are varied so that the value of $\tilde{\beta}$ is constantly equal to $1.02\times 10^{-1}\;$h-1. The dashed cyan line shows the speed of the infection wave obtained for Eq. (U2) with the same parameters. (b) Numerical solutions of Eq. (U3) compared with the position of the theoretical front (shown as a dashed red line). The parameters employed are the ones given in Table 1, with $\alpha=580\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1; we also set $R_{u}=L$ to focus on the infection inside an established tumour.
Let us first focus on the dynamics that do not present persistent, wide oscillations. The spatially homogeneous steady-state of the tumour is independent of $D_{v}$ and depends on $\alpha$ and $q$ only through $\tilde{\beta}$. This observation is already enough to provide useful insights on how to improve the overall treatment outcome, which remain valid whenever the infection spreads in the whole tumour. Indeed, increasing $\tilde{\beta}$ in any way and decreasing $q$ appears beneficial to the treatment. Even for parameter sets that maintain the equilibria unchanged, it is reasonable to expect differences in the spatial viral spread, which could significantly affect the treatment.
@@
The speed of travelling waves solutions for PDEs with spatial diffusion can be computed using linearisation techniques. In the case of Eq. (U2), infected cells invade regions in which uninfected cells are at carrying capacity at speed $2\sqrt{D(\beta-q)}$ and slow down as the uninfected cell density decreases [morselli23]; the same result can be also obtained in a rigorous way for a similar system of PDEs [dunbar84]. In the case of Eq. (U3), the minimal speed can be found at the intersection of two appropriate manifolds [baabdulla23]: although it is hard to derive an explicit expression, we can easily rely on numerical computations (see Appendix A). Fig. 1 shows the behaviour of the minimal speed for different parameter values. For the sake of simplicity, we restrict our attention to parameter combinations that maintain the ratio $\alpha/q_{v}$ constant: while this is not a biological constraint, it allows to focus our attention to the specific differences between the modelling approaches in comparable settings. As $q_{v}$ grows, the speed of the infection wave converges to the value $2\sqrt{D(\beta-q)}$, meaning that viral dynamics are irrelevant because the virus is cleared very quickly. Instead, the case $q_{v}=0$ corresponds to $\alpha=0$, which means that viral particles only diffuse without any production or decay; travelling waves are not possible in this hypotetical situation, even though this state has no biological meaning. It is also interesting to observe that for very low values of $q_{v}$ the speed is extremely low: in this case, viral dynamics are mostly dictated by spatial diffusion and the effects of the reaction part of the equation is negligible. In the intermediate cases, the speed is either higher or lower than $2\sqrt{D(\beta-q)}$ depending on whether $D_{v}$ is higher or lower than $D$. In the former case, it is interewsting to observe that there is a specific value of $q_{v}$ that maximises the speed.
@@
Let us now describe how the different speed of the infection affects the treatment in comparison to the previous approaches. We start by considering the lower values of $\alpha$ and $q_{v}$ from Table 1, as in this case we expect a more significant difference. Fig. 2, along with the video accompanying it (see electronic supplementary material S2), shows an excellent quantitative agreement between numerical solutions of the system of PDEs (U3) and the average over five numerical simulations of the corresponding agent-based model in one spatial dimension. Furthermore, the quasi-steady approximation of Eq. (2.3) holds most of the time, even in the agent-based model (despite some very small stochastic fluctuations), as it is evident from Fig. 2a. Nevertheless, the system’s transient behaviour significantly differs from the solution of Eq. (U2). Fig. 2a shows that the low viral decay allows the viral infection to reach the boundary of the tumour faster than in the case of cell-to-cell infection, in line with the higher speed predicted by theoretical results (see the vertical dash-dotted lines); as a consequence, uninfected cells start to decrease their numbers earlier (Fig. 2b). This difference does not significantly affect the system’s overall behaviour: in all the cases, the travelling wave of uninfected cells invades the surrounding area at the speed $2\sqrt{Dp}$ (vertical dashed lines in Fig. 2b-c) and cell densities at the centre of the tumour converge with damped oscillations to the equilibrium of the corresponding ODE (horizontal solid black lines in Fig. 2), given by Eq. (2.6).
@@
and it is shown as a yellow dashed line in Fig. 2b-c. In the full system, we observe a lower cell density at the front, which can be explained by the higher infiltration of the infection: the theoretical approximation (shown as a green dash-dotted line in Fig. 2b-c) predicts this decrease, but significantly underestimates the value; this disagreement is probably caused by the fact that the infection is not invading a spatially homogeneous uninfected population, which is the situation considered in the Appendix. Indeed, such a wave would have a longer spatial decay towards infinity, but this is not possible in the case of Fig. 2c because the uninfected population is more localised; as a consequence, the desired speed can only be obtained if the uninfected peak is slightly higher. In other words, nonlinear effects here play a more important role than in other settings.
Let us now consider higher values of $\alpha$ and $q_{v}$, keeping their ratio approximately as in the previous case. The electronic supplementary material S3 shows again an excellent quantitative agreement between numerical solutions of the system of PDEs (U3) and the average over five numerical simulations of the corresponding agent-based model in one spatial dimension. Furthermore, the initial invasion of infected cells shows a larger degree of similarity to the one obtained by the two-equation model in Eq. (U2) than in the previous case, as expected by the value of the invasion speed. After an initial transient time, the uninfected cell density is almost indistinguishable from the one resulting from Eq.  (U2). Travelling waves are remarkably similar to the solution of Eq.  (U2), with almost the same density of uninfected cells at the front: the theoretical value is a much better approximation than in the previous case and correctly predicts the front to be slightly lower than for the two equations model.
@@
Let us first consider the lower values of $\alpha$ and $q_{v}$. The behaviour of numerical simulations is very similar to the one observed in the previous case. Let us recall that, in the case of Eq. (P2), central infections remain localised in the initial position, as described in Ref. [morselli23] (see also Subsection 3.3); this is in sharp contrast with the outcome shown in Fig. 3. Indeed, in the present situation, viral particles can reach the outer region of the tumour without any obstacle and the constraints of cell movement do not stop the propagation of the infection. The similarity of this situation to the case of undirected cell movement suggests that it is more appropriate to compare it to Eq. (U2) than to Eq. (P2). Fig. 3a, along with the video accompanying it (see electronic supplementary material S2), shows that the spread of the infection is extremely similar with respect to the previous case, meaning that the viral diffusion affects the dynamics more significantly than cellular movement; small differences can only be observed by a close comparison of the PDE models, which however are totally unnoticeable in the agent-based model. The only relevant differences between undirected movement and standard diffusion appear to be the maximum value reached by uninfected cell density, which is higher in the former situation, and the profile of the final invasion front, which, in the latter case, is compactly supported and propagates to the surrounding tissues slightly faster. Overall, we can conclude that spatial viral dynamics have a more substantial effect on the evolution of the system than cellular spatial dynamics.
Fig. 3c, along with the video accompanying it (see electronic supplementary material S3), shows that also for high $\alpha$ and $q_{v}$ the outcome is very similar to the case of undirected cell movement, which confirms our previous observation that the infective dynamics are affected mainly by viral spatial spread. This is not trivial
@@ 3.3 Possible treatment failure: low viral diffusion
Lower viral diffusion clearly implies a slower invasion speed of infection, as also shown in Fig. 1. The movement of cells is still sufficient for spreading the infection all over the tumour, as shown in Fig. 4a-b. It is also interesting to observe that lower values of $\alpha$ and $q_{v}$ are now associated with a slower speed, differently from the case of faster viral diffusion: indeed, viral particles with a shorter lifespan cannot effectively propagate for long distances. Fig. 4a also shows that some stochastic events slow down the infection to the left of the domain, even though the results are averaged over five simulations; it appears therefore that stochasticity becomes more relevant for this less effective infection, although the overall behaviour is not affected significantly. Fig. 4b shows that for higher values of $\alpha$ and $q_{v}$ the behaviour is almost indistinguishable to the one of Eq. (U2): this is indeed the parameter range in which the quasi-steady assumption of the virus is most accurate. In both cases, the theoretical predictions of the peak of uninfected cells is very precise for the PDE and still constitutes a good approximation for the agent-based model.
@@
As the spatial spread of viral particles becomes less relevant, we expect to recover the main features of Eq. (P2). First of all, a central infection remains mostly confined at the centre of the tumour for both parameter sets that we consider (only shown in the electronic supplementary material S4): the central infection causes the total cell density to drop, while external uninfected cells proliferate; since cells cannot move towards an area with higher cell density and viral diffusion is negligible, the outer cells are never going to be infected and the tumour keeps expanding in the same way it would do in the absence of treatment, almost identically to the situation described in Ref. [morselli23].
We therefore focus on initial infections spread in the whole tumour, namely $R_{v}=R_{u}$, which gives rise to the formation of travelling waves for Eq. (P2) [morselli23]. For the sake of brevity, we restrict our attention to higher values of $\alpha$ and $q_{v}$, which facilitate the spread of the infection for the continuous model; we remark that both cases are shown in electronic supplementary material S5. Fig. 4c shows that the same kind of travelling waves is also obtained as a solution of Eq. (P3) and the two solutions superimpose almost exactly; the peak of uninfected cells is higher than the one of the solutions of Eqs. (U2) and (U3), in line with the less effective spread of the infection. Note that the agent-based model again shows that the infection is confined to the centre of the tumour, in line with the case that neglects explicit viral dynamics. It is important to stress that, for the current model, demographic stochasticity plays a much more important role than in previous models: any growth above the average of uninfected cells stops the movement of infected cells and hence cannot be compensated at later times by other processes.
@@
Figure 5: Numerical simulation of the discrete model with pressure-driven movement in two spatial dimensions at three different times. The parameters employed are the ones given in Table 1, with the exception of the diffusion coefficient of viral particles $D_{v}$ (which is set to $10^{-5}\;$mm2/h), the death rate of infected cells $q$ (which is set to $8.33\times 10^{-3}\;$h-1, i.e. one-fifth of the reference values) and the initial radius of viral infection $R_{v}$ (which is set equal to $R_{u}$); furthermore, we use the values $\alpha=3500\;$viruses/cells and $q_{v}=1\;$h-1. The dashed cyan circles in the top panels represent the expected positions of the uninfected invasion fronts in the absence of treatment, travelling at speed $\sqrt{D_{u}p/2}$. The dashed red circles in the bottom panels represent the front of the infected cells given by the numerical solution of Eq. (P3).
Reasonable increases of the effective infection rate $\tilde{\beta}$ do not lead to a better infiltration of the infection, as this causes a decrease in central cell density and creates the need for infected cells to move against a pressure gradient. Also, as we will show shortly, oscillations emerge in some parameter ranges associated with an effective infection, but are suppressed at low values of viral diffusion. On the other hand, building upon the results of Ref. [morselli23], we expect that decreasing the death rate of infected cells $q$ leads to a better outcome.